OpenStudy (osanseviero):

If cot=2/5, sen<0, how much is cos?

OpenStudy (osanseviero):

So adyacent is 2, and the hip is sqrt of 29, right?

OpenStudy (e.mccormick):

sen?

OpenStudy (osanseviero):

sin

OpenStudy (osanseviero):

sine*

OpenStudy (e.mccormick):

OK, if sine is negative, what quadrent is it in?

OpenStudy (osanseviero):

3 or 4, right?

OpenStudy (e.mccormick):

OK. So, it is in 3 or 4. Now, where could the cotangent be positive?

OpenStudy (osanseviero):

oh...in 4

OpenStudy (e.mccormick):

Cotangent positive in 4? Ummm....

OpenStudy (osanseviero):

cot is adjacent/opposite...so adjacent should be positive...right?

OpenStudy (osanseviero):

oh...but sine is negative...so the opposite is negative, so adjacent needs to be negative!

OpenStudy (e.mccormick):

This question seems to be using the unit circle definition. \(\cot\theta=\dfrac{\cos\theta}{\sin\theta}\) So for that to be positive, both the sine and cosine must have the same sign.

OpenStudy (e.mccormick):



OpenStudy (osanseviero):

Ok. everything is understood until there

OpenStudy (e.mccormick):

Have you done anything with the unit circle yet?

OpenStudy (e.mccormick):

"oh...but sine is negative...so the opposite is negative, so adjacent needs to be negative!" And that IS correct.

OpenStudy (e.mccormick):

See, there is a way to equate a right triangle to the unit circle... but not sure if you are using that. 

OpenStudy (osanseviero):

I have done it a long time ago, I know how to measure the trigonometric functions of a respective angle using the 45 triangles and the 90-60-30 triangles

OpenStudy (osanseviero):

Yep, we've done that before

OpenStudy (e.mccormick):

OK, so, what do you think the answer would have to be then?

OpenStudy (osanseviero):

Well, then we would need to measure the hypotenuse...sqrt of 29

OpenStudy (osanseviero):

and then cos should be 2/sqrt29 ...but that's not a choice

OpenStudy (osanseviero):

negative

OpenStudy (osanseviero):

oh...then I multiply by sqrt of 29/sqrt of 29 and I get the answer :D

OpenStudy (osanseviero):

Negative :P

OpenStudy (e.mccormick):

Ah, yah, you had to rationalize the denominator.

OpenStudy (osanseviero):

Yep :) forgot about that :P

OpenStudy (osanseviero):

Thanks a lot

OpenStudy (e.mccormick):

np. Have fun!